In a market for kitchen bags, the highest price consumers are willing to pay is $18 for a 48 pack and the lowest price producers are willing to accept is $8 per pack. The market equilibrium price is $10 per pack, at which 12 million packs are sold. (Assume that both demand and supply curves are straight lines.)

A.)In the market above, what is the consumer surplus ($ million)?

B.)In the market above, what are the total gains from trade ($ million)?
Answer #1

A. Consumer surplus is area below the demand curve and above the equilibrium price level.

Consumer surplus= 1/2*($18-$10)*12 million= $48 million

B. Total gains from trade is the total surplus. Total surplus is the area between the demand and the supply curve.

= 1/2*($18-$8)*12 million= $60 million